# A proof of Audenaert-Kittaneh’s Conjecture

### 10 Jan 2024 v1

@@ Abstract
We show that Audenaert-Kittaneh’s Conjecture related to $p$-Schatten class.
@@
Schatten $p$-norm, inequality
@@ 1 Introduction
The classical Clarkson’s inequalities C36 for the Lebesgue spaces $L_{p}$, and their non-commutative analogues for the Schatten trace ideals, play an important role in analysis, operator theory, and mathemetical physics. They have been generalised in various directions. Among these versions for more-general symmetric norms BH88 , for operators by $n$-tuples BK04 ; HK08 and for the Haagerup $L_{p}$ spaces FK86 , as well as refinements BCL94 . Let $\mathbb{B}(\mathscr{H})$ denote the algebra of all bounded linear operators acting on a complex separable Hilbert space $\mathscr{H}$. If $X\in\mathbb{B}(\mathscr{H})$ is compact, we denote by $\{s_{j}(X)\}$ the sequence of decreasingly ordered singular values of $X$. For $p>0$, let
@@
For $p=2$ both inequalities (1) and 2 reduce to the parallelogram law
@@
One natural generalizations of (1) and (2) are as follows:
@@
$n\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}.$ (6)
In view of the inequalities (3) and (4), it is reasonable to make the following conjecture concerning generalizations (along the lines of the inequalities (5) and (6) of the inequalities (1) and (2)) to $n$-tuples of operators.
Conjecture 1.2.
(AK12, , Audenaert & Kittaneh) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\right)^{\frac{q}{p}}$
@@
$n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$
@@
In CM16 , Conde and Moslehian obtained a weaker version of Conjecture 1.2 with a weaker coefficient $n^{\frac{q}{2}}$ instead of $n$, the conjecture 1.2 remains open.
@@
(CM16, , Theorem 3.2) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then
@@
for $p\in[2,\infty);\frac{1}{p}+\frac{1}{q}=1$,
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}$
for $p\in(1,2];\frac{1}{p}+\frac{1}{q}=1$.
@@ 2 Proof of Audenaert-Kittaneh’s Conjecture
In this section, we give a proof of Audenaert-Kittaneh’s Conjecture 1.2, the ideas are the same as those used in FK86 ; BK04 .
@@
$\left|f(z)\right|\leq n^{\frac{1}{2}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{2}}\left(\sum_{i=1}^{n}\|A_{i}(z)\|_{p}^{p}\right)^{\frac{1}{2}},$ (9)
@@
This proves the conjecture for $1<p\leq 2$. The reverse inequality for $2\leq p<\infty$ can be obtained from this by a duality argument.
@@ Acknowledgments
I heartily thank my advisor Minghua Lin for his guidance in my study and care in my life.

### 26 Feb 2024 v2

@@ Abstract
By using Hadamard’s 3-lines theorem for a certain analytic function defined in terms of the trace, we show that Audenaert-Kittaneh’s conjecture related to $p$-Schatten class.
@@
Schatten $p$-norm, inequality, Hadamard’s 3-lines theorem, trace function
journal: Journal of Functional Analysis
@@ 1 Introduction
The classical Clarkson’s inequalities C36 for the Lebesgue spaces $L_{p}$, and their non-commutative analogues for the Schatten trace ideals, play an important role in analysis, operator theory, and mathematical physics. They have been generalised in various directions. Among these versions for more-general symmetric norms BH88 , for operators by $n$-tuples BK04 ; HK08 and for the Haagerup $L_{p}$ spaces FK86 , as well as refinements BCL94 .
Let $\mathbb{B}(\mathscr{H})$ denote the algebra of all bounded linear operators acting on a complex separable Hilbert space $\mathscr{H}$. If $X\in\mathbb{B}(\mathscr{H})$ is compact, we denote by $\{s_{j}(X)\}$ the sequence of decreasingly ordered singular values of $X$. For $p>0$, let
@@
For $p=2$ both inequalities (1) and (2) reduce to the parallelogram law
@@
In BCL94 , Ball, Carlen and Lieb explained the following theorem fully.
Theorem 1.1 (Optimal 2-uniform convexity).
For $1\leq p\leq 2$, the inequality
$\left(\dfrac{\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}}{2}\right)^{\frac{2}{p}}\geq\|A\|_{p}^{2}+(p-1)\|B\|_{p}^{2}$ (3)
holds in the following case:
  1. 1.
    $A$ and $B$ are functions in $L_{p}$.
  2. 2.
    $A$ and $B$ are operators in $B_{p}(\mathscr{H})$.
If $2\leq p\leq\infty$, the inequality is reversed.
We note that validity of the inequality (3) implies 2-uniform convexity of $L_{p}$ and $C_{p}$. When $1\leq p\leq 2$, the function $f(t)=t^{\frac{p}{2}}$ is concave. So we can get the $\frac{p}{2}$-th power of the right side of (3)
$(\|A\|_{p}^{2}+(p-1)\|B\|_{p}^{2})^{\frac{p}{2}}$$\geq$ $2^{\frac{p}{2}-1}\|A\|_{p}^{p}+2^{\frac{p}{2}-1}(p-1)^{\frac{p}{2}}\|B\|_{p}^{p}$
$\geq$ $2^{\frac{p}{2}-1}(\|A\|_{p}^{p}+\|B\|_{p}^{p}).$
It implies that (3) is sharper than the first inequality of (1).
@@
One natural generalizations of partial (1) and (2) are as follows:
@@
$\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}.$ (7)
In view of the inequalities (4) and (5), it is reasonable to corresponding generalizations (along the lines of the inequalities (6) and (7) of the partial inequalities (1) and (2)) to $n$-tuples of operators. More precisely, in (AK12, , Section 8.1) entitled “Clarkson inequalities for several operators”, the authors presented the following conjecture:
Audenaert-Kittaneh’s Conjecture.
Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}},$ (8)
@@
$n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$ (9)
@@
The above inequalities Audenaert and Kittaneh conjectured is a crucial piece of the Clarkson-type inequality puzzle. Notice that both of (8) and (9) become equalities for $A_{1}=\ldots=A_{n}$, so these two inequalities are sharp. But little progress has been made in the decade since the conjecture was first proposed. It is worth mentioning that Conde and Moslehian obtained a weaker version of Audenaert-Kittaneh’s conjecture Audenaert-Kittaneh’s Conjecture with a weaker coefficient $n^{\frac{q}{2}}$ instead of $n$ in 2016.
@@
(CM16, , Theorem 3.2) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}},$
and for $p\geq 2$
@@
where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
@@ 2 Proof of Audenaert-Kittaneh’s Conjecture
In this section, we give a proof of Audenaert-Kittaneh’s conjecture Audenaert-Kittaneh’s Conjecture.
First, we introduced Hadamard three-lines theorem, which appears in the classic complex analysis textbook.
@@
Next, we give an elegant use of Hadamard’s three-lines lemma 2.4 for a certain analytic function defined in terms of the trace.
@@
$\left|f(z)\right|\leq n^{\frac{1}{2}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{2}}\left(\sum_{k=1}^{n}\|A_{k}\|_{p}^{p}\right)^{\frac{1}{2}},$ (12)
@@
This proves the conjecture for $1<p\leq 2$. The reverse inequality for $2\leq p<\infty$ can be obtained from this by a duality argument. ∎
Conflict of interest and data availability statement
The author reports no conflicting interests or funding. Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
@@ Acknowledgments
The author heartily thanks professor Eric A. Carlen for several valuable suggestions and comments, including introducing Hanner’s inequality and pointing out the result in this paper is best possible etc. The author would like to his advisor Minghua Lin for guidance in his study and care in his life.